OpenStudy (anonymous):

Making sure if I got the domain and range right - Identify the domain and range of the function: a = f(b) = 10b. - My answer in the comments.

OpenStudy (anonymous):

Domain: \[a \ge 0\] Range: \[10 \le b \ge 0\]

OpenStudy (anonymous):

Domain: Any number greater than or equal to zero. Range: Any number great than or equal to 10 or 0

OpenStudy (zzr0ck3r):

first off what you said does not make sense the way you wrote the function a=f(b) =10b impplies a is the output and b is the input, so domain would be in terms of b and range in terms of a your function is f(x) = 10x is there any reason we cant put in a negative number?

OpenStudy (anonymous):

The segment title says functions with rational numbers.

OpenStudy (zzr0ck3r):

I am not sure what you mean...

OpenStudy (anonymous):

The problem is from a problem book I have. It's divided into segments ie chapters. the chapter title says "Functions with Rational numbers."

OpenStudy (anonymous):

Oh and I forgot, it does say to exclude negative numbers on the chapter I'm on.

OpenStudy (anonymous):

a = 10b f(b) = 10b

OpenStudy (anonymous):

@Michele_Laino Sorry, I need to expertise again

OpenStudy (michele_laino):

ok! where is your question?

OpenStudy (anonymous):

Making sure if I got the domain and range right - Identify the domain and range of the function: a = f(b) = 10b.

OpenStudy (michele_laino):

is your function a real-value function?

OpenStudy (anonymous):

\[a \ge 0\]\[10 \le b \ge0\]

OpenStudy (anonymous):

Domain: Any number greater than or equal to zero. Range: Any number great than or equal to 10 or 0

OpenStudy (anonymous):

Negative numbers are excluded.

OpenStudy (michele_laino):

if \[0 \le b \le10\] then we can write: \[0 \le 10b \le 100\]

OpenStudy (anonymous):

b has to be 0 or 10 or greater than 10

OpenStudy (michele_laino):

so: \[0\le a=f(b) \le 100\]

OpenStudy (anonymous):

Range: a is great than 0, equal to or greater than 10?

OpenStudy (michele_laino):

if b is Greater or equal to zero the domain of your function is all positive real numbers, included zero, so the range of your function is the same set, namely all positive real numbers included zero graph of your function is: 

OpenStudy (anonymous):

Thanks, for the 3rd time :) Damn functions T.T

OpenStudy (michele_laino):

thanks!